# Compile Scene Graphs with Reinforcement Learning

### v1

@@ 4.2 How Well Do M-LLMs Reason About Visual Relationships?
Table 2: Comparison of VQA on the VG150 validation set across various models and settings. Gains compared to the Original Image (1st row) are indicated in red. “mask img.” refers to masking the entire image with random noise, “mask obj.” refers to masking object regions with black pixels, “wo cats.” refers to not providing object categories in the prompt, and “wo box.” refers to not providing bounding boxes in the prompt.
@@
org. img. 2.26 1.94 45.75 45.61 28.69 29.17 53.74 53.35
mask img. 1.00 (-1.26) 1.04 (-0.90) 21.80 (-23.95) 21.61 (-24.00) 3.85 (-24.84) 3.95 (-25.22) 0.03 (-53.71) 0.02 (-53.33)
mask obj. 1.89 (-0.37) 1.86 (-0.08) 37.15 (-8.60) 37.16 (-8.45) 12.79 (-15.90) 13.22 (-15.95) 16.23 (-37.51) 16.82 (-36.53)
wo cats. 2.50 (+0.24) 2.38 (+0.44) 32.83 (-12.92) 32.68 (-12.93) 9.46 (-19.23) 10.11 (-19.06) 16.78 (-36.96) 18.06 (-35.29)
   + mask img. 0.97 (-1.29) 1.00 (-0.94) 15.41 (-30.34) 15.25 (-30.36) 0.01 (-28.68) 0.01 (-29.16) 0.18 (-53.56) 0.22 (-53.13)
   + mask obj. 1.75 (-0.51) 1.65 (-0.29) 27.91 (-17.84) 28.37 (-17.24) 3.26 (-25.43) 3.77 (-25.40) 4.67 (-49.07) 5.52 (-47.83)
wo box. 26.01 (+23.75) 25.93 (+23.99) 61.93 (+16.18) 61.32 (+15.71) 53.46 (+24.77) 52.10 (+22.93) 78.13 (+24.39) 77.10 (+23.75)
   + mask img. 10.14 (+7.88) 10.19 (+8.25) 36.26 (-9.49) 35.23 (-10.38) 11.47 (-17.22) 11.44 (-17.73) 0.01 (-53.73) 0.00 (-53.35)
   + mask obj. 19.26 (+17.00) 19.08 (+17.14) 54.23 (+8.48) 53.84 (+8.23) 33.46 (+4.77) 33.23 (+4.06) 40.26 (-13.48) 39.26 (-14.09)
We evaluate the visual relationship reasoning capabilities of open-source multimodal LLMs using a four-to-one Visual Question Answering (VQA) task. Each model is prompted with an image and a corresponding question. The used prompt template is:
Analyze the relationship between the object "{sub_name}" at {sub_box} and the object "{obj_name}" at {obj_box} in an image of size ({width}x{height}). The bounding boxes are in [x1, y1, x2, y2] format. Choose the most appropriate relationship from the following options: A) {choices[0]}; B) {choices[1]}; C) {choices[2]}; D) {choices[3]}.
We report Acc (accuracy over all questions) and mAcc (mean accuracy per image) in Table 2. The results reveal that many multimodal LLMs struggle with visual relationship reasoning. Moreover, the task exhibits a noticeable text bias, and the presence of bounding boxes can sometimes mislead the model’s attention. As a simpler task compared to SGG, the poor performance suggests that directly applying multimodal LLMs to SGG may yield suboptimal results.
Table 3: SGDET Performance of Qwen2VL-7B/2B-Instruct models on the VG150 validation set.
Method Params Failure Rate (%) AP50 R@20 R@50
w.o. predefined classes or predicates
baseline 7B 47.26 6.75 0.84 0.84
zero-RL 7B 0.08 (-47.18) 13.97 (+7.22) 13.11 (+12.27) 13.11 (+12.27)
SFT 7B 40.16 13.66 9.15 9.15
SFT+RL 7B 0.04 (-40.12) 17.52 (+3.86) 20.16 (+11.01) 20.18 (+11.03)
[…]
zero-RL 2B 0.06 (-69.62) 6.19 (+4.46) 4.34 (+4.25) 4.34 (+4.25)
SFT 2B 75.56 7.23 4.80 4.80
SFT+RL 2B 0.12 (-75.44) 16.58 (+9.35) 18.70 (13.90) 18.70 (13.90)
w. predefined classes & predicates
baseline 7B 55.32 6.03 0.76 0.76
zero-RL 7B 0.08 (-55.24) 11.63 (+5.60) 10.67 (+9.91) 10.67 (+9.91)
SFT 7B 38.50 14.58 9.92 9.96
SFT+RL 7B 0.10 (-38.40) 18.72 (+4.14) 20.40 (+10.48) 20.47 (+10.51)
[…]
zero-RL 2B 0.08 (-56.48) 8.03 (+5.72) 6.13 (+5.96) 6.13 (+5.96)
SFT 2B 73.80 7.81 5.13 5.13
SFT+RL 2B 0.28 (-73.52) 17.20 (+9.39) 18.82 (+13.69) 18.87 (+13.74)
(a) Failure Rate vs Training Steps
(b) AP@50 vs Training Steps
(c) R@50 vs Training Steps
(d) Failure Rate vs Training Steps (2B)
(e) AP@50 vs Training Steps (2B)
(f) R@50 vs Training Steps (2B)
(g) Failure Rate vs Training Steps (7B)
(h) AP@50 vs Training Steps (7B)
(i) R@50 vs Training Steps (7B)
Figure 2: Performance of Qwen2VL-2B/7B-Instruct during SFT / RL over training steps. For SFT, the model was trained with a batch size of 128 for 3 epochs. For RL, the model was trained for 1 epoch with a batch size of 32, and the group size of GRPO was set to 8. w.o. cats. refers to the setting without predefined object classes or predicate categories.
4.3 How Well do M-LLMs Generate Scene Graphs?
The results demonstrate that reinforcement learning (RL) significantly reduces the failure rate and enhances both object detection and relationship recognition. In contrast, supervised fine-tuning (SFT) alone results in a relatively high failure rate and limited performance improvements. As shown in Fig. 2, the failure rate rapidly declines to a very low level when using RL, whereas SFT continues to suffer from a high failure rate. Meanwhile, our experimental results suggest that predefined object classes or relationship categories are unnecessary, despite their potential to reduce the search space of M-LLMs.
5 Conclusion
This work explores the application of reinforcement learning (RL) to multimodal large language models (LLMs) for enhancing end-to-end scene graph generation (SGG). From the perspective of scene graphs, we design rule-based rewards, including node-level, edge-level, and format-based rewards. These reward formulations contribute to stable and effective policy optimization for scene graph tasks. We open-source our framework to foster further research and development in advancing multimodal LLMs for visual reasoning.

### v3

@@ 4.2 How Well Do M-LLMs Reason About Visual Relationships?
We evaluate the visual relationship reasoning capabilities of open-source multimodal LLMs using a four-to-one Visual Question Answering (VQA) task. Each model is prompted with an image and a corresponding question. The used prompt template is: Analyze the relationship between the object "{sub_name}" at {sub_box} and the object "{obj_name}" at {obj_box} in an image of size ({width}x{height}). The bounding boxes are in [x1, y1, x2, y2] format. Choose the most appropriate relationship from the following options: A) {choices[0]}; B) {choices[1]}; C) {choices[2]}; D) {choices[3]}. We report Acc (accuracy over all questions) and mAcc (mean accuracy per image) in Table 7. The results reveal that many multimodal LLMs struggle with visual relationship reasoning. Moreover, the task exhibits a noticeable text bias, and the presence of bounding boxes can sometimes mislead the model’s attention. As a simpler task compared to SGG, the poor performance suggests that directly applying multimodal LLMs to SGG may yield suboptimal results.
Table 1: SGDET performance on the VG150 validation set. For M-LLMs, predefined object classes and relation categories are included in the input prompts.
Method Params Failure Rate (%) AP@50 Recall mRecall
Specific Models
IMP [34] - - 20.91 17.85 2.66
MOTIFS [38] 29.56 27.21 7.84
VCTree [29] 28.13 24.87 8.47
OvSGTR [4] 33.39 26.74 5.83
Commercial M-LLMs
GPT-4o [9] - 2.94 0.00 0.00 0.00
Gemini 1.5 Flash [25] - 1.10 0.51 0.10 0.08
Gemini 2.0 Flash [6] - 1.06 0.54 0.07 0.03
Open-sourced M-LLMs
LLaVA v1.5 [21] 7B 82.70 0.00 0.00 0.00
Qwen2-VL-2B-Instruct [32] 2B 59.96 2.18 0.07 0.18
   +SFT 2B 72.42 8.10 5.47 1.46
Qwen2-VL-7B-Instruct [32] 7B 54.46 6.07 0.69 0.80
   +SFT 7B 39.54 14.18 9.62 3.30
[…]
R1-SGG 2B 0.10 17.87 21.09 7.48
R1-SGG-Zero 7B 0.04 15.59 18.34 8.32
R1-SGG 7B 0.08 19.47 23.75 11.43
4.3 How Well do M-LLMs Generate Scene Graphs?
4.3.1 Benchmark on VG150
  • Specific Models: Models built upon specific detectors such as Faster R-CNN [26] (e.g., IMP [34], MOTIFS [38]) or DETR [1] (e.g., OvSGTR [4]) for scene graph generation.
  • Commercial M-LLMs: Advanced multimodal LLMs such as GPT-4o [9] and Gemini 1.5 Flash [25].
  • Open-source M-LLMs: Publicly available models such as LLaVA v1.5 [21], Qwen2-VL [32], and our proposed R1-SGG-Zero and R1-SGG.
    • R1-SGG-Zero: Based on Qwen2-VL-2B/7B-Instruct, fine-tuned with GRPO but without supervised fine-tuning (SFT).
    • R1-SGG: Built on the same backbone, fine-tuned with GRPO and initialized from SFT checkpoints.
The results in Table 1 reveal several key observations.
Zero-shot Performance of M-LLMs. Either commercial or open-source multimodal LLMs struggle to generate accurate scene graphs, and this can be attributed to several factors. First, the internal processing of private models such as GPT-4o remains opaque to users, resulting in suboptimal object detection performance. Second, models like LLaVA v1.5 align visual and textual features only at the image level, typically using a fixed resolution of 336$\times$336, which restricts spatial understanding. Third, although models such as Gemini 2.0 and Qwen2-VL demonstrate a degree of spatial understanding, the task of scene graph generation is much complex than pure object detection or visual grounding. Consequently, their zero-shot performance drops significantly.
SFT vs. RL. 1) RL substantially improves performance across all metrics compared to SFT alone. Specifically, RL dramatically reduces the failure rate (e.g., from 72.42% to 0.10% for 2B models) and yields significant gains in AP@50, Recall, and mRecall. This highlights the effectiveness of GRPO in enhancing the model’s ability to generate accurate and complete scene graphs. 2) SFT achieves moderate improvements in AP@50 and Recall over the baseline but struggles with a relatively high failure rate. This suggests that SFT primarily improves relation prediction while being less effective at correcting structural errors, such as missing objects, relationships, or format inconsistencies. 3) applying RL on top of SFT (i.e., R1-SGG) further boosts performance over both SFT and R1-SGG-Zero in most cases. This indicates that combining SFT and RL benefits from better initialization, leading to stronger relation recognition and higher recall. 4) larger models (e.g., 7B) consistently outperform smaller models (e.g., 2B) across AP@50, Recall, and mRecall, demonstrating the benefits of scaling model capacity for scene graph generation.
Compared to Specific Models. The gap between AP@50 and Recall highlights the advantage of dense predictions. However, our models, such as R1-SGG, achieve a notable mean Recall (mRecall) of 11.43%, suggesting that multimodal LLMs are more effective at generating less biased scene graphs. Moreover, specific models are typically restricted to a limited vocabulary and struggle to generalize across domains, whereas multimodal LLMs exhibit greater adaptability and broader generalization capabilities.
Overall, the results demonstrate that reinforcement learning (RL) significantly reduces the failure rate and enhances both object detection and relationship recognition. In contrast, supervised fine-tuning (SFT) alone results in a relatively high failure rate and limited improvements. As shown in Fig. 2, the failure rate quickly drops to near-zero with RL, whereas SFT continues to suffer from frequent structural errors.
4.3.2 Benchmark on PSG
Table 2: Performance on the PSG dataset [36]. For M-LLMs, predefined object classes and relation categories are included in the input prompts.
Model Params Failure Rate (%) AP@50 Recall mRecall
Specific Models
IMP [34] 16.50 6.50
MOTIFS [38] 20.00 9.10
VCTree [29] - - - 20.60 9.70
GPSNet [19] 17.80 7.00
PSGFormer [36] 18.60 16.70
Open-sourced M-LLMs
LLaVA v1.5 [21] 7B 81.97 0.07 0.00 0.00
TextPSG [42] - - - 4.80 -
ASMv2 [33] 13B 0.87 21.45 14.77 11.82
LLaVA-SpaceSGG [35] 13B - - 15.43 13.23
Qwen2-VL-2B-Instruct 2B 67.20 4.89 0.39 0.26
   +SFT 2B 6.54 36.05 22.06 14.92
Qwen2-VL-7B-Instruct 7B 37.97 12.75 3.18 4.33
   +SFT 7B 0.96 40.79 24.73 17.11
[…]
R1-SGG 2B 2.70 39.28 38.49 31.21
R1-SGG-Zero 7B 0.00 32.92 37.00 32.04
R1-SGG 7B 0.00 42.05 43.48 33.71
As shown in Table 2, our R1-SGG approach achieves strong performance on the PSG dataset. Compared to baselines, SFT significantly improves AP@50, Recall, and mean Recall (mRecall), while reinforcement learning further enhances relationship recognition, achieving the highest Recall (43.48% for 7B model) and mRecall (33.71%). Notably, our method also drives the failure rate to zero, demonstrating the effectiveness of reinforcement learning in promoting structured, accurate scene graph generation even without predefined object categories.
4.4 Qualitative Results
We present qualitative results in Fig. 4. In this example, the ground-truth scene graph (Fig. 4-(a)) captures key objects and their relationships but is biased toward the predicate “has”. In contrast, the zero-shot Qwen2-VL-7B-Instruct (Fig. 4-(b)) fails to generate a valid JSON output, indicating poor instruction-following ability. With supervised fine-tuning (SFT), the model produces structurally valid graphs (Fig. 4-(c)) but frequently omits important relationships, resulting in a sparse scene graph. R1-SGG-Zero (7B), trained with RL only, improves relational recall and structure (Fig. 4-(d)), yet still outputs inaccurate triplets such as $<$wheel, on, horse$>$ and $<$helmet.2, on, horse$>$. Finally, R1-SGG (7B), trained with both SFT and RL, produces a complete and structurally consistent scene graph (Fig. 4-(e)), with results that even surpass the ground truth in relational richness.
4.5 Discussion
Through the exploration of applying GRPO to the SGG task, we make several observations.
KL Regularization. We compare models trained with and without KL divergence regularization in Fig. 3. From the result, removing KL regularization leads to improved performance, particularly with a significant reduction in failure rate.
Sampling Length. In our experiments, the default sampling length is set to 1,024, which sufficiently covers most corrected answers. As shown in Fig. 3, increasing the sampling length to 2,048 does not yield further performance improvements, suggesting that longer sampling might enlarge the search space and introduce additional optimization difficulties without clear benefits. This observation aligns with prior findings on test-time scaling, where increasing Chain-of-Thought (CoT) length can degrade performance [39].
Group Size. As shown in Fig. 3, increasing the group size from 8 to 16 stabilizes training performance, consistent with the intuition that more candidates reduce variance in group statistics estimation. To balance computational cost and performance, we adopt a group size of 8 as the default in this work.
To Think or Not Think? We adopt the <think>$\cdots$</think><answer>$\cdots$</answer> format in the system prompt, following DeepSeek R1 [8]. However, models such as Qwen2-VL-2B/7B-Instruct often fail to produce outputs with the <think> tag after fine-tuning, indicating difficulty in adhering to the intended structure. This suggests that rule-based rewards alone are insufficient to trigger abstract reasoning patterns like CoT, and highlights the need for additional SFT on CoT-specific datasets to incentivize coherent intermediate reasoning.
Generalization Across Datasets.
Table 3: Generalization across datasets using Qwen2-VL-7B-Instruct as the baseline. Columns under Pre-training indicate whether the weights were initialized from specific checkpoints, while the Training column specifies the dataset(s) used during the fine-tuning stage. “w/o cats.” denotes prompts without predefined object classes or relation categories.
Model Pre-Training Training VG150 PSG
Failure Rate AP@50 Recall mRecall Failure Rate AP@50 Recall mRecall
baseline - - 54.46 6.07 0.69 0.80 37.97 12.75 3.18 4.33
baseline (w/o cats.) - - 44.58 6.83 0.61 0.37 30.28 13.79 1.96 2.30
SFT - VG150 39.54 14.18 9.62 3.30 22.10 11.05 3.03 1.36
SFT (w/o cats.) - VG150 42.98 13.03 8.94 2.47 19.81 12.15 3.87 1.81
R1-SGG-Zero - VG150 0.04 15.59 18.34 8.32 0.18 24.92 13.83 8.90
R1-SGG-Zero (w/o cats.) - VG150 0.06 15.30 16.33 6.94 0.18 18.10 6.16 3.38
R1-SGG SFT VG150 0.08 19.47 23.75 11.43 0.23 18.12 9.10 5.13
R1-SGG (w/o cats.) SFT (w/o cats.) VG150 0.30 18.09 22.73 9.62 0.64 14.64 7.51 3.88
[…]
SFT (w/o cats.) - PSG 2.54 7.94 1.77 1.25 1.01 39.02 23.70 17.17
R1-SGG-Zero - PSG 0.12 14.22 8.90 5.34 0.00 32.92 37.00 32.04
R1-SGG-Zero (w/o cats.) - PSG 0.02 9.08 2.80 1.78 0.05 24.26 19.94 18.04
R1-SGG SFT PSG 0.94 10.38 4.40 2.69 0.00 42.05 43.48 33.71
R1-SGG (w/o cats.) SFT (w/o cats.) PSG 0.14 9.38 2.16 1.55 0.14 41.15 41.44 31.51
We report performance comparisons across datasets, including VG150 and PSG, in Table 3 to evaluate the generalization ability of models trained with SFT and RL. The results reveal several key observations: 1) VG150 poses a significantly greater challenge than PSG. For instance, SFT trained solely on PSG achieves a high AP@50 of 40.58 and Recall of 24.75%, with a low failure rate of 0.91%. In contrast, SFT trained only on VG150 results in a much higher failure rate of 39.54%, with notably lower AP@50 (14.18) and Recall (9.62%). 2) SFT has a strong domain-specific effect. SFT models trained on one dataset (e.g., VG150) exhibit substantial performance drops when evaluated on another (e.g., PSG), reflecting limited transferability. For example, VG150-trained SFT only achieves 3.03% Recall and 1.36% mRecall on PSG. 3) Initialization of RL matters. R1-SGG initialized with SFT checkpoints consistently outperforms R1-SGG-Zero. On VG150, R1-SGG (7B) achieves 23.75% Recall and 11.43% mRecall versus 18.34% and 8.32% for R1-SGG-Zero. A similar trend is observed on PSG. This highlights the importance of using SFT as a warm-start for reinforcement learning, which leads to improved sample efficiency and stronger downstream performance. 4) R1-SGG-Zero exhibits stronger cross-dataset generalization. This aligns with the domain-specific nature of SFT—models trained via SFT tend to overfit to the source domain, resulting in degraded performance on unseen datasets. In contrast, R1-SGG-Zero, trained without SFT, generalizes more robustly across domains.
Table 4: Ablation study of reward formulations on the VG150 validation set. All results are obtained with the R1-SGG (7B).
Setting Sparsity Metric Aligned Failure Rate (%) AP@50 Recall (%) mRecall (%)
Hard Recall sparse 0.08 19.47 23.75 11.43
Hard Recall + Relax medium 0.02 19.93 24.05 9.61
Soft Recall dense 0.06 18.73 21.92 5.61
Hard Recall vs. Soft Recall. As shown in Table 4, Hard Recall outperforms other variants despite providing sparser reward signals. This highlights the importance of aligning reward functions with evaluation metrics, rather than prioritizing reward smoothness alone.
5 Conclusion
We present a reinforcement learning framework for enhancing end-to-end Scene Graph Generation (SGG) with multimodal large language models (M-LLMs). To align training with the structured nature of scene graphs, we design a set of rule-based rewards, comprising three scene graph variants (Hard Recall, Hard Recall+Relax, and Soft Recall) and a format consistency reward, which enable fine-grained and stable policy optimization via GRPO. Our approach significantly improves the structural validity and relational accuracy of generated scene graphs. We release our code and models to support future research on structured visual understanding with M-LLMs.
Appendix A Supplementary Material
A.1 Prompt Templates for SGG
In this work, we adopt two prompt templates for scene graph generation, as illustrated in Table 6 and Table 5. The difference lies in whether predefined object classes and relation categories are provided.
Table 5: Prompting an M-LLM to generate scene graphs without providing predefined object classes or predicate types.
messages = [{ "role": "system", "content": " {system_prompt}" }, { "role": "user", "content": f"""Generate a structured scene graph for an image using the following format: “‘json { "objects": [ {"id": "object_name.number", "bbox": [x1, y1, x2, y2]}, … ], "relationships": [ {"subject": "object_name.number", "predicate": "relationship_type", "object": "object_name.number"}, … ] }“‘. ### **Guidelines:** - **Objects:** - Assign a unique ID for each object using the format "object_name.number" (e.g., "person.1", "bike.2"). - Provide its bounding box ‘[x1, y1, x2, y2]’ in integer pixel format. - Include all visible objects, even if they have no relationships. - **Relationships:** - Represent interactions accurately using "subject", "predicate", and "object". - Omit relationships for orphan objects. ### **Example Output:** “‘json { "objects": [ {"id": "person.1", "bbox": [120, 200, 350, 700]}, {"id": "bike.2", "bbox": [100, 600, 400, 800]}, {"id": "helmet.3", "bbox": [150, 150, 280, 240]}, {"id": "tree.4", "bbox": [500, 100, 750, 700]} ], "relationships": [ {"subject": "person.1", "predicate": "riding", "object": "bike.2"}, {"subject": "person.1", "predicate": "wearing", "object": "helmet.3"} ] } “‘ Now, generate the complete scene graph for the provided image: """ } ]
Table 6: Prompting an M-LLM to generate scene graphs with predefined object classes and predicate types. Here, OBJ_CLS and REL_CLS refer to the predefined object classes and relation categories respectively.
messages = [{ "role": "system", "content": " {system_prompt}" }, { "role": "user", "content": f"""Generate a structured scene graph for an image using the following format: “‘json { "objects": [ {"id": "object_name.number", "bbox": [x1, y1, x2, y2]}, … ], "relationships": [ {"subject": "object_name.number", "predicate": "relationship_type", "object": "object_name.number"}, … ] }“‘. ### **Guidelines:** - **Objects:** - Assign a unique ID for each object using the format "object_name.number" (e.g., "person.1", "bike.2"). The **object_name** must belong to the predefined object set: ‘{OBJ_CLS}’. - Provide its bounding box ‘[x1, y1, x2, y2]’ in integer pixel format. - Include all visible objects, even if they have no relationships. - **Relationships:** - Represent interactions accurately using "subject", "predicate", and "object". - Omit relationships for orphan objects. - The **predicate** must belong to the predefined relationship set: ‘{REL_CLS}’. ### **Example Output:** “‘json { "objects": [ {"id": "person.1", "bbox": [120, 200, 350, 700]}, {"id": "bike.2", "bbox": [100, 600, 400, 800]}, {"id": "helmet.3", "bbox": [150, 150, 280, 240]}, {"id": "tree.4", "bbox": [500, 100, 750, 700]} ], "relationships": [ {"subject": "person.1", "predicate": "riding", "object": "bike.2"}, {"subject": "person.1", "predicate": "wearing", "object": "helmet.3"} ] } “‘ Now, generate the complete scene graph for the provided image: """ } ]
A.2 How Well Do M-LLMs Reason About Visual Relationships?
@@
Table 7: Comparison of VQA on the VG150 validation set across various models and settings. Gains compared to the Original Image (1st row) are indicated in red. “mask img.” refers to masking the entire image with random noise, “mask obj.” refers to masking object regions with black pixels, “w/o cats.” refers to not providing object categories in the prompt, and “w/o box.” refers to not providing bounding boxes in the prompt.
@@
org. img. 2.3 1.9 45.8 45.6 28.7 29.2 53.7 53.4
mask img. 1.0 (-1.3) 1.0 (-0.9) 21.8 (-24.0) 21.6 (-24.0) 3.9 (-24.8) 4.0 (-25.2) 0.0 (-53.7) 0.0 (-53.4)
mask obj. 1.9 (-0.4) 1.9 (-0.1) 37.2 (-8.6) 37.2 (-8.4) 12.8 (-15.9) 13.2 (-16.0) 16.2 (-37.5) 16.8 (-36.5)
w/o cats. 2.5 (+0.2) 2.4 (+0.4) 32.8 (-12.9) 32.7 (-12.9) 9.5 (-19.2) 10.1 (-19.1) 16.8 (-36.9) 18.1 (-35.3)
   + mask img. 1.0 (-1.3) 1.0 (-0.9) 15.4 (-30.3) 15.3 (-30.3) 0.0 (-28.7) 0.0 (-29.2) 0.2 (-53.6) 0.2 (-53.1)
   + mask obj. 1.8 (-0.5) 1.7 (-0.3) 27.9 (-17.8) 28.4 (-17.2) 3.3 (-25.4) 3.8 (-25.4) 4.7 (-49.1) 5.5 (-47.8)
w/o box. 26.0 (+23.7) 25.9 (+24.0) 61.9 (+16.2) 61.3 (+15.7) 53.5 (+24.8) 52.1 (+22.9) 78.1 (+24.4) 77.1 (+23.8)
   + mask img. 10.1 (+7.9) 10.2 (+8.2) 36.3 (-9.5) 35.2 (-10.4) 11.5 (-17.2) 11.4 (-17.7) 0.0 (-53.7) 0.0 (-53.4)
   + mask obj. 19.3 (+17.0) 19.1 (+17.1) 54.2 (+8.5) 53.8 (+8.2) 33.5 (+4.8) 33.2 (+4.1) 40.3 (-13.4) 39.3 (-14.1)
To evaluate the reasoning capabilities of M-LLMs over visual relationships, we present results in Table 7. We vary both the visual input and the text prompt conditions to assess robustness. For visual variations, we consider: org. img., mask img., and mask obj.; for prompt variations, we add: w/o cats. (without object categories) and w/o box. (without bounding boxes).
A.3 Discussion
004004004004008008008008001,20012001{,}2001 , 20020202020404040406060606080808080100100100100Training StepsFailure Rate (%)004004004004008008008008001,20012001{,}2001 , 2000055551010101015151515Training StepsAP@50004004004004008008008008001,20012001{,}2001 , 20000555510101010Training StepsRecall (%)004004004004008008008008001,20012001{,}2001 , 2001,60016001{,}6001 , 6000050505050100100100100Training StepsFailure Rate (%)004004004004008008008008001,20012001{,}2001 , 2001,60016001{,}6001 , 600005555101010101515151520202020Training StepsAP@50004004004004008008008008001,20012001{,}2001 , 2001,60016001{,}6001 , 600001010101020202020Training StepsRecall (%)
Qwen2-VL-2B-Instruct (SFT) Qwen2-VL-7B-Instruct (SFT)
R1-SGG-Zero (2B) R1-SGG-Zero (7B) R1-SGG (2B) R1-SGG (7B)
Figure 2: Comparison of R1-SGG-Zero and R1-SGG models against SFT baselines (Qwen2-VL-2B/7B-Instruct) across training steps on the VG150 validation set in terms of Failure Rate (%), AP@50, and Recall (%).
004004004004008008008008001,20012001{,}2001 , 2001,60016001{,}6001 , 6000020202020404040406060606080808080Training StepsFailure Rate (%)004004004004008008008008001,20012001{,}2001 , 2001,60016001{,}6001 , 6005555101010101515151520202020Training StepsAP@50004004004004008008008008001,20012001{,}2001 , 2001,60016001{,}6001 , 6005555101010101515151520202020Training StepsRecall (%)004004004004008008008008001,20012001{,}2001 , 2001,60016001{,}6001 , 6000050505050100100100100Training StepsFailure Rate (%)004004004004008008008008001,20012001{,}2001 , 2001,60016001{,}6001 , 600005555101010101515151520202020Training StepsAP@50004004004004008008008008001,20012001{,}2001 , 2001,60016001{,}6001 , 600001010101020202020Training StepsRecall (%)004004004004008008008008001,20012001{,}2001 , 2001,60016001{,}6001 , 6000050505050100100100100Training StepsFailure Rate (%)004004004004008008008008001,20012001{,}2001 , 2001,60016001{,}6001 , 6005555101010101515151520202020Training StepsAP@50004004004004008008008008001,20012001{,}2001 , 2001,60016001{,}6001 , 6005555101010101515151520202020Training StepsRecall (%)
  R1-SGG (2B, w. KL)   R1-SGG (2B, w/o KL)
  R1-SGG (2B, sampling length 2k)   R1-SGG (2B, sampling length 1k)
  R1-SGG (2B, group size 16)   R1-SGG (2B, group size 8)
Figure 3: Performance comparison of R1-SGG (2B) across training steps on the VG150 validation set. Each row evaluates a different setting: (Top) KL divergence regularization ($\beta{=}0.04$ vs. $\beta{=}0$), (Middle) sampling length, and (Bottom) group size. Metrics reported include Failure Rate (%), AP@50, and Recall (%).
Refer to caption Refer to caption (a) GT.Refer to caption Refer to caption (b) Qwen2-VL-7B-Instruct.Refer to caption Refer to caption (c) Qwen2-VL-7B-Instruct (SFT).Refer to caption Refer to caption (d) R1-SGG-Zero (7B).Refer to caption Refer to caption (e) R1-SGG (7B).
Figure 4: Qualitative comparison of generated scene graphs. (a) Ground-truth scene graph annotated by humans. (b) Zero-shot Qwen2-VL-7B-Instruct produces an invalid JSON (failure to follow format). (c) Qwen2-VL-7B-Instruct (SFT) outputs a valid graph but omits many relations. (d) R1-SGG-Zero (7B) trained with RL only recovers most objects and relations but still hallucinates incorrect triplets (e.g., $<$ wheel, on, horse$>$ and $<$ helmet.2, on, horse $>$). (e) R1-SGG (7B) fine-tuned with SFT + RL yields a complete, structurally correct scene graph with higher recall.